OpenStudy (anonymous):

What would be solution set f(x)=7x^3-5x^2-63x+45

jhonyy9 (jhonyy9):

x^2(7x-5)-9(7x-5)=(x^2-9)(7x-5)=(x-3)(x+3)(7x-5)

OpenStudy (anonymous):

{3,-3,5/7}